OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 120r3 square inches. Each tile has an area of sqrt(16r^9). The total number of tiles used can be represented by the expression below.

OpenStudy (anonymous):

@reemii @robz8 @Jhannybean ???

OpenStudy (jhannybean):

@jim_thompson5910 :)

OpenStudy (jhannybean):



jimthompson5910 (jim_thompson5910):

Total Area = (# of tiles)*(Area per tile) 120r^3 = (x)*(sqrt(16r^9)) solve for x and rationalize the denominator if necessary

OpenStudy (jhannybean):

\[\large 120r^3= x*\sqrt{16r^9}\] simplify \[\large \sqrt{16r^9} \to \sqrt{16} * \sqrt{r^9}\] first... which is \[\large \sqrt{16} = \sqrt{2*2*2*2}\] and \[\large \sqrt{r^9} = r^{9/2}\]

OpenStudy (anonymous):

I'm on the final question of my Apex math. I need help. Thank you.

OpenStudy (jhannybean):

ok? But this isn't even your question to be posting that on.

OpenStudy (jhannybean):

After you simplify that, @kyleguitarplayer we can proceed